OpenStudy (anonymous):

PLEASE HELP!!!! A pyramid is placed inside a prism as shown. The pyramid has the same base area, B, as the prism but half the height, h, of the prism. Which expression gives the volume of the pyramid? Its not v=1/3*b*h my choices v=2/3bh v=1/4bh v=2bh v=1/6bh

OpenStudy (superdavesuper):

"half the height, h, of the prism" <- that is why "Its not v=1/3*b*h" u should be able to pick the right choice from here :)

OpenStudy (lucaz):

the volume of a pyramid is (1/3)*b*h, half of the prism's height = (1/2)*h volume would be (1/3)*b*(1/2)*h = (1/6)bh

OpenStudy (anonymous):

@superdavesuper thank so its V=16bh

OpenStudy (superdavesuper):

missing a / there ;)

OpenStudy (anonymous):

typo v=1/6bh

OpenStudy (superdavesuper):

bingo @binks

OpenStudy (anonymous):

:) thanks

OpenStudy (superdavesuper):

welcome